half life period of a radioactive element is 140 days. after 560 days 1 g of the element will reduce to ......

Grade:

3 Answers

vikas askiitian expert
13 years ago

A/Ao = (1/2)t/T     

A is final amount  , t is time , & T is half life ...

 

A = Ao(1/2)t/T

A = Ao(1/2)560/140

A = Ao(1/2)4 = (1/16)Ao

Ao = initial amount = 1g

A = 1/16gm = 0.0625gms

this is the amount of  element after 560 days

om harne
9 years ago
A wave undergoes reflection from a rigid boundary. One of its characteristic parameters that changes is
(a) frequency
(b) phase
(c) velocity
(d) wavelength
sugandha
9 years ago
phase